OpenStudy (anonymous):

@ganeshie8 Do you know how to do this? Write the vector v in terms of i and j whose magnitude and direction angle θ are given. ||v||= 8, θ = 30°

ganeshie8 (ganeshie8):

@ikram002p u free ? plz help..

OpenStudy (ikram002p):

no he dnt ^^

OpenStudy (ikram002p):

and no im not free :)

ganeshie8 (ganeshie8):

lol thats funny :) ok let me show u bquick how to do this

ganeshie8 (ganeshie8):

wat you're given is simply \((r, \theta) = (8, 30)\)

ganeshie8 (ganeshie8):

\(i\) and \(j\) are the x and y components of that polar representation

OpenStudy (anonymous):

x = r cos(theta) y = r sin(theta)

ganeshie8 (ganeshie8):

so, \(i = x =r \cos \theta = 8*\cos 30 = ?\) \(j = y =r \sin \theta = 8*\sin 30 = ?\)

OpenStudy (anonymous):

i=1.23? j=-7.90?

ganeshie8 (ganeshie8):

i=8*cos(30) = 8*sqrt(3)/2 = 4sqrt(2)

ganeshie8 (ganeshie8):

j=8*sin(30) = 8*1/2 = 4

ganeshie8 (ganeshie8):